OpenStudy (anonymous):

Solve the equation: 2*4^x+4^-x=3 (Hint: Multiply by something)

OpenStudy (anonymous):

take log on both sides then log 2^a=a log 2 use this to solve

OpenStudy (anonymous):

assume 4^x =a then the q becomes 2a+1/a= 3 or 2a^2-3a+1=0 2a^2-2a -a+1=0 or 2a(a-1) -1(a-1) =0 or (2a-1)(a-1)=0 hence a=1 or a=1/2 if a=1 then 4^x =1 means x=0 if a=1/2 then 4^x =1/2 means x=-1/2

OpenStudy (anonymous):

Thank you matricked!!!!!!!! Helped alot!!

OpenStudy (anonymous):

welcome ...